OpenStudy (anonymous):

Consider two straight lines L1 (y = 2x +1) and L2 (y=x). a) What is the slope of line L1 in respect to line L2? b) What is the equation of line L1 in respect to line L2?

OpenStudy (anonymous):

Oh, you want the angle between the two lines.

OpenStudy (anonymous):

You need to use the formula to find the angle between two lines. \[\tan \alpha = \frac{m_2-m_1}{1+m_1m_2}\]

OpenStudy (anonymous):

m_2 is typically the greater slope, which is 2 in this case. m_1 is therefore 1.

OpenStudy (anonymous):

Since the slope of a line is defined as the tangent of that line with respect to the x-axis, the tangent of the angle between your two lines should give you the relative slope...which I assume is what the question wants.

OpenStudy (anonymous):

hmm... ok, I will try it again with this extra knowledge

OpenStudy (anonymous):

right, i think i get it

OpenStudy (anonymous):

I could be misinterpreting your question here - it's late where I am and I'm about to hit the hay.

OpenStudy (anonymous):

BecomeMyFan, did you get the second part? I forgot, sorry. I'm about to go to bed. Hopefully you can lure someone in to finish it...

OpenStudy (anonymous):

full problem is Consider two straight lines L1 (y = 2x +1) and L2 (y=x). a) What is the slope of line L1 in respect to line L2? b) What is the equation of line L1 in respect to line L2? NOTE. Think that the original xy co-ordinate system has been rotated in respect to the origin so that line L2 defines the new x-axis x´ and y´ is the new y-axis!